# Mixed Diffusion for 3D Indoor Scene Synthesis

### v1

@@ E.2 Unconditional 3D Scene Synthesis
DiffuScene 61.12 0.46 53.70 $\pm$ 2.94 1.24 45.04 0.70 51.53 $\pm$ 1.05 1.10 43.30 0.21 53.50 $\pm$ 2.84 0.67
MiDiffusion (ours) 65.72 2.83 54.93 $\pm$ 3.85 1.12 43.61 0.42 51.21 $\pm$ 1.33 0.71 43.66 0.84 53.16 $\pm$ 2.32 0.61
@@
DiffuScene-[41] 17.21 0.70 52.15 0.35 32.60 0.72 55.50 0.22 36.18 0.88 57.81 0.21
DiffuScene 17.43 0.82 51.09 $\pm$ 0.47 0.66 33.07 0.93 53.82 $\pm$ 4.01 0.34 35.27 0.58 54.24 $\pm$ 3.65 0.36
MiDiffusion (Ours) 21.32 2.62 54.93 $\pm$ 3.85 0.32 33.05 0.87 52.32 $\pm$ 2.34 0.16 36.21 1.43 54.86 $\pm$ 2.59 0.16
Table 9: Evaluation results for unconditional 3D scene synthesis against training scenes and using layouts rendered on 3.1m squares.
@@ Appendix F Additional Layout Examples
F.1 Qualitative Examples for Additional Applications
We show qualitative examples of additional applications of our models as explained in Sec. 5.2.3. Fig. 11 and 11 are two examples of label constrained scene synthesis. In particular, Fig. 11 is conditioned on five dining chairs, which occur less often then even number of chairs. Our model is still able to generate reasonable layout.
We show label constrained scene completion examples in Fig. 13 and Fig. 13. With our corruption-and-masking strategy, our models are able to complete scene layout under object category constraints.
Figure 10: Label constrained bedroom scene synthesis.
Figure 11: Label constrained dining room scene synthesis.
Figure 12: Label constrained bedroom scene completion.
Figure 13: Label constrained living room scene completion.
F.2 Additional Layouts for Floor-Conditioned Scene Synthesis
We provide additional layout image generated by MiDiffusion in our floor-conditioned scene synthesis experiment. These layouts are generated in a sequence of random sampling of test floor plans after removing floor plan duplicates.

### v2

@@ E.4 Unconditional 3D Scene Synthesis
DiffuScene [46] 61.12 0.46 51.09 $\pm$ 0.47 1.24 45.04 0.70 51.53 $\pm$ 1.05 1.10 43.30 0.21 53.50 $\pm$ 2.84 0.67
MiDiffusion (ours) 62.02 1.12 52.36 $\pm$ 1.54 0.95 43.37 0.68 51.21 $\pm$ 0.82 0.70 42.43 0.77 53.46 $\pm$ 1.93 0.58
@@
DiffuScene (from [46]) 17.21 0.70 52.15 0.35 32.60 0.72 55.50 0.22 36.18 0.88 57.81 0.21
DiffuScene [46] 17.43 0.82 51.09 $\pm$ 0.47 0.66 33.07 0.93 53.82 $\pm$ 4.01 0.34 35.27 0.58 54.24 $\pm$ 3.65 0.36
MiDiffusion (Ours) 18.17 0.82 52.36 $\pm$ 1.54 0.16 31.39 1.15 52.38 $\pm$ 1.85 0.18 32.49 1.20 51.86 $\pm$ 1.80 0.15
Table 11: Evaluation results for unconditional 3D scene synthesis against training scenes and using layouts projected to 6.2 squares.
DiffuScene was originally proposed for 3D scene synthesis without floor conditioning. Although this is not the main focus of this paper, we provide experimental results comparing MiDiffusion and DiffuScene for completeness using the released model weights. Since this is an unconditioned problem, we slightly modify the proposed architecture in Fig. 1 by removing modules related to the conditional input, including the floor plan feature encoder, the second AdaLN and the second multi-head cross-attention unit in each transformer block. We re-use all other hyper-parameters from the floor-conditioned experiments.
We follow the same layout image generation and evaluation method as in Sec. 5, except we project rendered scenes to a white background without floor plans. The results are presented in Table 10. We also note that in [46], DiffuScene used a 6.2m square to capture all synthesized layouts when generating the orthographic projection images for evaluation. Since the dining room and living room training data contains objects beyond this range, some of the predicted objects will be outside the projection range leading to incomplete result images for evaluation. In addition, the synthesized results were compared to the ground-truth training split. For completeness, we follow the evaluation setup in DiffuScene and present quantitative results in Table 11 using the same predicted layouts as in Table 10. The first row is directly copied from [46] demonstrating that we reproduced DiffuScene results close to the original publication values. The only exception is CA%, for which DiffuScene does not release their classification network. Therefore, we adopt the AlexNet [25] based classifier from ATISS [32] to be consistent with Sec. 5.
Both Table 10 and  11 show that MiDiffusion achieves similar results, despite that its transformer-based denoising network is chosen for floor-conditioned scene synthesis.
@@ Appendix F Additional Layout Examples
We provide additional layout images generated by MiDiffusion in our floor-conditioned scene synthesis experiment. These layouts are generated in a sequence of random sampling of test floor plans after removing floor plan duplicates.